IVVICII Ulslal ILS A toroidal solenoid has 560 turns, cross-sectional area 6.90 cm, and mean radius 4.10 cm. Part A CalcualtPart C The current is directed from terminal a of the coil to terminal b. Is the direction of the induced emf from a to b orPlease solve for A,B and C
Answer #1

(a) Coils sey Inductance - » N- 560 turns A = 6.90 x 10+ m² = 0.0 41m =) Lets me sex inductance eq; A = N o = N (B. A) = N(6) Induced End; Er? - TOAS Induced rug ΛΙ 2.00 5.00 at 0,003 & E= - L di at = - (10.55 X 10-4) (-10%) – 10.55 X 10-1 2 1.05